OpenStudy (anonymous):

4x+4y= 2x^2, solve for x

OpenStudy (anonymous):

4y=2x^2+4x

OpenStudy (anonymous):

what does x =?

OpenStudy (anonymous):

start with \[2x^2-4x-4y=0\] then \[x^2-2x-2y=0\] then either use the quadratic formula or complete the square

OpenStudy (anonymous):

\[x^2-2x=2y\] \[(x-1)^2=2y+1\] \[x-1=\pm\sqrt{2y+1}\] \[x=1\pm\sqrt{2y+1}\] is the complete the square method

OpenStudy (anonymous):

Then do you use the FOIL method?

OpenStudy (anonymous):

for quadratic equation?